Starting with her next salary payment (next month), Kathy intends to save $200 each month. if the interest rate is 19 per month, how much will Kathy have saved after 12 months (12 deposits altogether over 12 months)? Fill in the blank with the correct answer, rounded to 2 decimal places. 12 months later (in month 12), Kathy would havein the bank account
Answer #1

So at the end of 12th month kathy will receive $225.36

Solution -

for 12 month she will receive (9months) = 202*(1+0.01)^12= 225.36

Formula -

A = P*(1+i)^n

A= amount

P= Principal = 200

i= interest rate = 1% monthly

n= number of periods = 12 months